OpenStudy (anonymous):

Plz help quick will medal and fan! Evaluate. 8*3+6 ------- -2[6/2-(9-1)] =______

OpenStudy (anonymous):

Plz help Quick!

OpenStudy (er.mohd.amir):

8*3+6=24+6=30

OpenStudy (anonymous):

can you say that again?

OpenStudy (er.mohd.amir):

-2[(6/2-(9-1)]=-2[3-8)=-2*-5=10

RhondaSommer (rhondasommer):

8*3+6-2[6/2-(9-1)] is the equation right?